OpenStudy (anonymous):

Red team scores 3 runs in one inning, then in the next inning, they score 1, then they score zero, and they continue this pattern. Blue team scores 2 points every next inning, how many innings will it take for Team Red and Team Blue to have the same score.

OpenStudy (anonymous):

3 2 1 = 6 2 2 2 = 6 3 Innings

OpenStudy (anonymous):

but red team would score 3 1 0, 3 1 0, not 321

OpenStudy (gw2011):

If I understand the problem correctly, this is what the pattern should be for R (Red) and B (Blue): R: 3 1 0 3 1 0 3 1 0 B: 2 2 2 2 2 2 2 2 2 Therefore, after 2 inninngs both teams have 4 points each (R=3+1=4 and B=2+2=4)